OpenStudy (anonymous):

what is the trigonometric identity for ( half angle) sin^2x= EMERGENCY last thing I need to know before my test.. Please help guys.. Well appreciative

OpenStudy (shane_b):

Google :) http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

im not allowed to open that up.. theres a lock on my computer.. can u help me shane.

OpenStudy (shane_b):

The lock isn't working well if you're allowed here then :P \[sin^2 \theta=\frac{1}{2}(1=cos(2 \theta))\]

OpenStudy (shane_b):

Typo in there...that should have been: \[sin^2 \theta=\frac{1}{2}(1-cos(2 \theta))\]

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

and lol it just prevents me from opening stuff up